OpenStudy (anonymous):

1/8>1/11. explAin how you know this is true. use this number line

OpenStudy (anonymous):



OpenStudy (helder_edwin):

u have \[ \large \frac{1}{8}\ ?\ \frac{1}{11} \] if u cross multiply u get \[ \large 1\cdot11\ ?\ 1\cdot8 \]

OpenStudy (helder_edwin):

11>8 so \[ \large \frac{1}{8}>\frac{1}{11} \]

OpenStudy (anonymous):

ohh ok

OpenStudy (helder_edwin):

for istance. which is greater 2/5 or 7/8?

OpenStudy (helder_edwin):

\[ \large \frac{2}{5}\ ?\ \frac{7}{8} \]

OpenStudy (anonymous):

7/8 ?

OpenStudy (helder_edwin):

yes because \[ \large 2\cdot8<7\cdot5 \]

OpenStudy (anonymous):

ooohh ok